A cup of beans cost ₦8, how much would a similar tin, three times the height and diameter, full of beans cost.
Answer by VFBundy(438)

The area for each the tins is: h*¶r².

Let's say the first tin has a height of 1 and a diameter of 1. This means the total area is: h*¶r² = (1)*¶(0.5)². (Remember, the radius (r) is half of the diameter.) Anyway, the total area of the first tin is:

(1)*¶(0.5)² = 1 * 0.25¶ = 0.25¶.

If the bigger tin has a height and diameter that are each three times bigger, the height is 3 and the diameter is 3. So., the total area of the second tin is:

h*¶r² = (3)*¶(1.5)² = 3 * 2.25¶ = 6.75¶.

So, the smaller tin has a total area of 0.25¶ and the larger tin has a total area of 6.75¶. This means the bigger tin is 27 times bigger. (Since 6.75¶/0.25¶ = 27.) Since the larger tin is 27 times bigger, that means the cost of the beans will be 27 times larger than the beans in the smaller tin. That means the cost of the beans in the larger tin is: ₦8 * 27 = ₦216.